North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size

Q6 - 2/3 November 2021

Ian Pearson reports a cool fine day with some bubbly clouds at Grosmont on Tuesday with only Ian and Bill Dobson to start with, although Bill could not stay for long due to other commitments with a horse. As Bill was leaving, Gordon Wells arrived. Then around lunchtime Jon Bradley clocked on, so Ian wasn’t Billy no mates!

After the PAT testing from last Thursday with Nigel Bill in attendance, there were some electrical repairs to carry out to which Gordon attended, but leaving the oil filled radiator to repair next time, as it needs a new cable and plug. Ian put all the tested equipment that had passed, back in the workshop in their respective places. The Defibrillator, which was checked out last Thursday and all was OK, started giving a warning bleep in the morning, and showing a red light indicating the battery was ready for replacement. Nigel Bill was informed and Richard Wheeler is ordering a replacement. The replacement stove chimney extension had been fitted sometime over the weekend by Andy Whiffin, but, as it was raining when he was on site, he could not repair the cracks in the two roofing sheets which were leaking. He will be back later to deal with them.

Jon spent all afternoon removing some of the old fence and fitting a new gate to give a safer access to Deviation Shed. This critical path can now be used to give access from the MPD, although there is still some finishing off to do which Jon will complete shortly. The old access, direct on to the ash pit running line, will be kept to allow heavy items to be moved into Deviation Shed when required, but will be kept locked when not needed.

Wednesday, in contrast, was a day of rain and hail showers, some of which were very heavy. Six were present to brave the rain - Bill Dobson, Steve Hyman, Ed Bolam, Nigel Hall, David Potter and Ian Pearson himself.

It turned into a joint J27/Q6 working party and on the Q6, Bill and David removed the whistle valve for examination and gave it to Owain for machining. They then turned their attention to the lefthand big end, which was found to be a lot warmer than the righthand side during last week's running. The large cotter was slightly eased and a new securing cotter made and fitted. Ian fitted 16 new trimmings to the horn guide brass oil boxes.

With the operating season now over until Christmas, and both locomotives only recently returned to traffic, it has been decided to reduce our working parties to one day a week for the moment. So, the next working party will be on Wednesday 10 November.